OpenStudy (anonymous):

How do i work this? y= x+4 -3 + 2y = 8

OpenStudy (aripotta):

substitute x + 4 for y in the second equation

OpenStudy (aripotta):

so -3 + 2(x + 4) = 8

OpenStudy (aripotta):

then use the distributive property, combine like terms, and solve for x

OpenStudy (anonymous):

so you would do -3+2 which = -1? then the rest is distrubitive property?

OpenStudy (aripotta):

no, distributive property first. distribute the 2 into the parentheses

OpenStudy (anonymous):

oh so 2x+8 then what lol sorry i suck at math.

OpenStudy (aripotta):

so then our equation would be -3 + 2x + 8 = 8 add like terms

OpenStudy (anonymous):

@ForevernAlways you should check that what you wrote in your question is actually the problem you are working on.

OpenStudy (anonymous):

mmm i would you put that? bc u alredi put +8 but then you added =8....im confused Dx

OpenStudy (aripotta):

this was our equation to begin with: -3 + 2(x + 4) = 8. after distributing the 2, we get -3 + 2x + 8 = 8

OpenStudy (anonymous):

okay lol so after you get that, you combine -3 and 8?

OpenStudy (aripotta):

yep

OpenStudy (aripotta):

are you good now? or do you need help still?

OpenStudy (anonymous):

nope im good :3 thank you so much <3 can i finish this, and when i get the awnser can i ask u if its right lol. i cant fail this stupid homework thingy

OpenStudy (anonymous):

too late.

OpenStudy (aripotta):

sure, that wouldn't be a problem

OpenStudy (aripotta):

>_>

OpenStudy (anonymous):

okai hehe thanks <3

OpenStudy (anonymous):

@Algebraic! what are you talking about?

OpenStudy (aripotta):

i think she was implying that you've already failed it...

OpenStudy (aripotta):

but i won't let that happen

OpenStudy (aripotta):

just ask me for help

OpenStudy (anonymous):

oh thats mean! but thanks hun (:

OpenStudy (aripotta):

i thought it was mean too. and help is always given at Hogwarts to those who deserve it

OpenStudy (anonymous):

is x=-1.5 right

OpenStudy (anonymous):

lol i love hogwarts <3

OpenStudy (aripotta):

it would just be 1.5. no negative

OpenStudy (aripotta):

do you have to find out what y equals too?

OpenStudy (anonymous):

oh okay thanks for the help <3 oh dang yeah i do, how do i do that. Sorry im re taking algebra on the computer and its so much harder then regular at skwl and plus im takin geometry lol

OpenStudy (aripotta):

all you have to do is add 1.5 and 4. since we know that y = x + 4 and x = 1.5

OpenStudy (anonymous):

ohhh yay harldy no work hehe.

OpenStudy (aripotta):

yep :) all the work went into finding out what x was lol

OpenStudy (anonymous):

yep xD